OpenStudy (wampominater):

verify the trigonometric equation by substituting identities to match the right hand side with the left hand side: (sinx / 1-cosx) + (sinx / 1+cosx) = 2cscx

OpenStudy (wampominater):

looks like this \[\frac{ sinx }{ 1-cosx } + \frac{ sinx }{ 1+cosx }=2cscx\] been stuck on this forever...

OpenStudy (maheshmeghwal9):

on left hand side take L.C.M.

OpenStudy (butterflydreamer):

start with your left hand side :) hint: You'll want put the two fractions over the same denominator .

OpenStudy (wampominater):

so, would i get them both under 1-cos^2x?

OpenStudy (wampominater):

over*

OpenStudy (butterflydreamer):

correct:)

OpenStudy (maheshmeghwal9):

It Becomes \[\frac{ sinx(1+cosx)+sinx(1-cosx) }{ \sin^2x }\]

OpenStudy (wampominater):

alright then do I distribute sin?

OpenStudy (butterflydreamer):

yyeeepp

OpenStudy (butterflydreamer):

distribute sinx :)

OpenStudy (wampominater):

\[\frac{ 2sinx }{ \sin^2x }\] like this?

OpenStudy (maheshmeghwal9):

yes

OpenStudy (butterflydreamer):

then from there, you can simplify the fraction and you'll get the answer (RHS) :)

OpenStudy (wampominater):

Ok so how do I simplify this, Sorry! this is where i have been stuck at.

OpenStudy (butterflydreamer):

oh you can divide the numerator and denominator by sinx

OpenStudy (wampominater):

so is it 2/sinx? And then just change it to csc?

OpenStudy (butterflydreamer):

likeee: